solve log base 5(x+2) + log base 8(x+4) = log abse 8(2) + log base 8(4)
Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

solve log base 8(x+2) + log base 8(x+4) = log abse 8(2) + log base 8(4)
----------
log8[(x+2)(x+4)] = log8[2*4]
------
x^2+6x+8 = 8
------
x^2+6x = 0
x(x+6) = 0
x = 0 is the only acceptable answer
------
Cheers,
Stan H.
-------------

Answer by MathTherapy(10837)

solve log base 5(x+2) + log base 8(x+4) = log abse 8(2) + log base 8(4)

Does this start with: log5(x + 2) or  log8(x + 2)? 

The person who responded has the latter. Does he know something I don't?